# Local asymptotics of selection models with applications in Bayesian selective inference

### v2

@@ 3.2 Inference on winners
where $Z_{n}^{1}=\sqrt{n}(\bar{Y}_{n}-\mu)$ and $Z_{n}^{2}=\sqrt{n}(V_{n}-\sigma^{2}+(\bar{Y}_{n}-\mu)^{2})$. The selection function is $p_{n}(y^{n})=\mathbf{1}(s>t)$, where $s$ is the $t$-value as defined above (using only the first $n_{1}$ observations). The selection function of the asymptotic Gaussian model on $Z_{n}$ is $p_{n}^{*}(z;\theta)=\mathbb{E}_{\theta}[p_{n}(Y^{n})\mid Z_{n}=z]$, which does not admit, to the best of our knowledge, a simple analytical expression. Despite the apparent simplicity of the model, derivation of the appropriate asymptotic model for inference is complex, as the asymptotic selection function is the conditional acceptance probability of a subsample $t$-test given the full-sample sufficient statistics $(\bar{Y}_{n},V_{n})$. This, in turn, induces a selection probability function $\varphi_{n}^{*}(h)$ on a Gaussian location model $Z\sim N(h,I_{\theta}^{-1})$ which is non-standard: it cannot be represented by a standard selection criterion on the Gaussian model, such as involving a significance test on $h$. Figure 2 shows one-dimensional cuts of $p^{*}_{n}(z;\theta)$, $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r_{n}^{*}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ for $n_{1}=20$, $n_{2}=20$, $\theta=(0,1)$ and $t=1$. Under the true parameter, the selection probability is $\varphi_{n}(0,1)\approx 0.16$. The first two functions were computed via numerical integration, and $r_{n}^{*}(h;\theta)$ via Monte Carlo integration of $p^{*}_{n}(z;\theta)$. As expected, the approximation of the true selection probability by a Gaussian probability is much more accurate in the direction of the mean ($h_{1}$) than in the direction of the variance ($h_{2}$): in this problem $Z_{n}^{1}$ is exactly normally distributed. Additionally, since the selection condition is employed to determine large values of $\mu$, the selection function fluctuates much more in the first coordinate, corresponding to $\mu$ (top left plot), than in the second one (top right).
@@ 4 Asymptotic posterior distributions
Below we elaborate further on the distinction between these posteriors.
@@
The growing popularity of selection-adjusted posteriors underscores the importance of understanding their theoretical properties, particularly as their behaviour deviates from that of standard Bayesian models. This section makes two primary contributions. We prove that the consequences of Theorem 1 extend to the Bayesian framework, leading to a result analogous to the Bernstein von-Mises Theorem for the context of selection. Specifically, we show that the asymptotic shape of the posterior distribution matches that of the corresponding Gaussian selection model under a uniform prior. These latter models provide a more intuitive framework that can be leveraged to better understand the former. Furthermore, as a consequence of the previous result, we indicate a key practical limitation of Bayesian selective inference: probabilistic claims about the parameter (e.g. that it lies within a certain credible interval with probability 90%), cannot be given a frequentist interpretation with high accuracy. In order to achieve frequentist-matching properties, more complex prior distributions that depend on the sample size are required.
@@ 4.1 Asymptotic behavior of selective posteriors
$E_{\theta_{0}}(T_{n})\to 0\quad\text{and}\quad\sup_{\|\theta-\theta_{0}\|\geq\varepsilon}E_{\theta_{0}}(1-T_{n})\to 0.$ (16)
@@
where $Z_{n}=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})\rvert_{\theta_{0}}$. That is, the Total Variation distance between the true and asymptotic posteriors vanishes asymptotically in probability conditionally on selection.
@@
Figure 3 compares the posterior densities for the local parameter $h=\sqrt{n}(\theta-\theta_{0})$, $\pi_{h,n}(h\mid y^{n})$, for two exponential models with their theoretical Gaussian limits. We examine two of the scenarios described at the end of Section 3: deterministic selection and randomization with Gaussian noise. The model is $Y_{1},\ldots,Y_{50}\sim\text{Exp}(\theta)$, $\theta\sim\text{Gamma}(1,0.1)$, and the true parameter is $\theta_{0}=2$. All posterior densities correspond to an observed sample average of $\bar{y}_{n}=0.45$. The true posterior densities based on the exponential likelihood, $\pi_{h,n}(h\mid y^{n})$, are shown in blue, while the Gaussian approximations $\pi^{*}_{h,n}(h\mid z_{n})$, as predicted by the theorem, are shown in black. For comparison, the posterior density in the absence of selection, which is approximately a rescaled $N(1/\bar{y}_{n},\theta_{0}^{2}/n)$ under standard theory, is shown in red. Two selection mechanisms are considered: a deterministic one (left), with selection condition $\bar{Y}_{n}<0.5$, and a randomized one (right), with selection condition $\bar{Y}_{n}+0.2\times N(0,1/n)<0.5$. In both cases, the selection probabilities $\varphi_{n}(\theta)$ and $\varphi^{*}_{n}(h)$ were obtained via numerical integration, as in the numerical investigation presented in Figure 1. We can readily observe that both posterior densities are non-Gaussian, exhibit higher uncertainty than the non-selective ones, have slowly decaying tails in the direction of the parameter space where $\varphi^{*}_{n}(\theta)$ vanishes, and demonstrate marked skewness, particularly in the case involving deterministic selection.
@@
For further numerical illustration, we revisit the example discussed in Section 3.2, inference on a normal mean, under a selection condition defined by a $t-$ test on a subsample. We assume true parameter value $\theta_{0}=(0,1)$ and consider sample sizes $n_{1}=100,n_{2}=50$, with threshold $t_{n}=1$, which corresponds to a selection probability of about 0.15. Figure 4 illustrates, for a particular data sample, the marginal posteriors for $h_{1}$ and $h_{2}$, derived from the exact posterior $\pi_{h,n}(h\mid y^{n})$, constructed assuming the prior $\pi(\mu,\sigma^{2})\propto 1/\sigma$, and the Gaussian approximation $\pi^{*}_{h,n}(h\mid z_{n})$. In each case, the full posterior for the local parameter $h=(h_{1},h_{2})$ is constructed by MCMC sampling. The vertical lines in the figure indicate the limits of 90% credible intervals for $h_{1},h_{2}$ constructed from the exact posterior $\pi_{h,n}(h\mid y^{n})$. In the case of $h_{1}$, for instance, this interval is $(-2.11,2.50)$ and has probability content under the marginal posterior corresponding to the Gaussian approximation of 0.905. The corresponding figure in the case of $h_{2}$ is 0.884. The marginal posteriors in the true selection model and the Gaussian approximation model match closely, but not exactly.
@@
$(20,20)$ 0 0.898 (0.051) 0.850 (0.075)
1 0.905 (0.047) 0.869 (0.063)
2 0.916 (0.039) 0.906 (0.054)
3 0.914 (0.048) 0.924 (0.064)
$(50,25)$ 0 0.907 (0.036) 0.871 (0.055)
1 0.917 (0.031) 0.888 (0.047)
2 0.913 (0.054) 0.903 (0.048)
3 0.892 (0.099) 0.917 (0.051)
$(100,50)$ 0 0.908 (0.027) 0.884 (0.040)
1 0.915 (0.029) 0.892 (0.037)
2 0.918 (0.045) 0.905 (0.034)
3 0.886 (0.130) 0.909 (0.039)
4.2 Frequentist calibration of selection-adjusted Bayesian inference
@@
In the classical IID framework, the Bernstein-von Mises Theorem establishes asymptotic equivalence between the Bayesian and frequentist approaches to inference on $\theta$. In practical terms, this implies that Bayesian posterior probabilities can be interpreted as frequentist probabilities under the assumption that there is a fixed, true data-generating parameter $\theta_{0}$. For instance, $1-\alpha$ Bayesian credible intervals are also asymptotic $1-\alpha$ frequentist confidence intervals. This equivalence stems from the fact that Gaussian location models with a uniform prior satisfy this probability-matching condition exactly.
@@
In the presence of selection, this is no longer the case: selection-adjusted posterior distributions can remain significantly miscalibrated as the sample size increases. This follows by combining Proposition 1 with the observation that the uniform prior on the location parameter lacks a Bayesian-frequentist equivalence under selection (Proposition 2). Achieving approximate frequentist calibration in this setting typically requires the use of more carefully tuned priors that account for the selection mechanism and vary with the sample size: see García Rasines and Young (2022). This is not surprising, since, typically in selective inference, the dependence of the selection function on $n$ does not vanish asymptotically. This consideration of mis-calibration is relevant both for Bayesians and for frequentist practitioners who use Bayesian methods for their computational or interpretive convenience.
@@
In general, given a one-dimensional parameter of interest $\psi=g(\theta)$, a prior density for $\theta$ is probability-matching if $\Pi(g(\theta_{0})\mid Y)\sim U(0,1)$ under repeated sampling of $Y$ when $\theta_{0}$ is the true parameter, where $\Pi(\cdot\mid Y)$ is the marginal posterior distribution function of $\psi$ induced by the said prior. This ensures that credible intervals are valid confidence intervals, as
$P_{\theta_{0}}\{\Pi^{-1}(\alpha_{1}\mid Y)\leq\psi_{0}\leq\Pi^{-1}(\alpha_{2}\mid Y)\}=P_{\theta_{0}}\{\alpha_{1}\leq\Pi(\psi_{0}\mid Y)\leq\alpha_{2}\}=\alpha_{2}-\alpha_{1},$
@@
for all $\theta_{0}$ and $10<\alpha_{1}<\alpha_{2}<$ 1. Conveniently, improper uniform priors on location parameters are probability-matching. Thus, by the Bernstein-von Mises Theorem, any fixed prior density which is strictly positive at $\theta_{0}$ is first-order probability matching, meaning that the distribution of $\Pi(\psi_{0}\mid Y)$ converges to a $U(0,1)$ at a rate of $1/\sqrt{n}$, and hence credible intervals are asymptotically valid confidence regions. The following proposition shows that, in wide generality, in a selective Gaussian model, unlike the non-selective counterpart, there is no fixed prior which is, even asymptotically, probability matching. This includes the uniform prior $\pi(\theta)\propto 1$. Consequently, the limiting Gaussian posteriors indicated in Corollary 1 are not calibrated in a frequentist sense.
@@
Let $Y\sim N(\theta,\Sigma)\in\mathbb{R}^{p}$, with $\theta\in\mathbb{R}^{p}$ unknown and $\Sigma$ known and positive definite. Let $p(y)$ be a selection function such that
$\sum_{j=1}^{p}\text{Cov}(Y_{i},Y_{j})\frac{\partial}{\partial\theta_{j}}\varphi(\theta)\neq 0\quad\text{for all }\theta\in\mathbb{R}^{p}.$
Then, no prior density is probability-matching for $\theta_{i}$ in the selective model for $Y$.
This proposition relates to a well-known result due to Lindley (1958), which states that the only exact probability-matching priors in one-dimensional models are uniform priors for location parameters. In this case, the condition assumed that $\varphi(\theta)$ is not constant along the lines $\{[\text{Cov}(Y_{i},Y_{1}),\ldots,\text{Cov}(Y_{i},Y_{p})]t\colon t\in\mathbb{R}\}$ prevents the selective model from being a location model in these directions. The proof proceeds by finding a transformation of the original parameter $\theta\to(\theta_{i},\chi)$ such that $\theta_{i}$ and $\chi$ are orthogonal, and imposing that $\varphi$ not be constant in $\theta_{i}$ in this parametrization. Although seemingly complex, this condition ought to be satisfied by any reasonable selection rule which is used to identify interesting values of the parameter of interest $\theta_{i}$. If $Y_{i}$ is uncorrelated with the remaining variables, the required condition is simply
$\frac{\partial}{\partial\theta_{i}}\varphi(\theta)\neq 0.$
The degree of departure from uniformity of the marginal posterior density will generally depend on how strong selection bias is, that is how much the selection mechanism alters the distribution of the data under the true parameter, measured through the partial derivatives of $\varphi(\theta)$.
The lack of frequentist calibration can be made more precise in the one-dimensional setting, when the selection function of the Gaussian model is monotonic. In such cases, the posterior distribution tends to overstate, under repeated sampling, regions of the parameter space with low selection probability, i.e. it overcompensates selection bias. The following result extends Proposition 1 in García Rasines and Young (2022).
@@
$P_{\theta_{0}}\{\theta_{0}\leq\Pi^{-1}(\alpha\mid Y)\mid u\}<\alpha\quad\forall(\alpha,\theta_{0})\in(0,1)\times\mathbb{R}.$ (18)
The proof of this result hinges on the existence of a function $\pi(\theta;y)$ which is strictly increasing in $\theta$ for every $y$ such that, if employed as a data-dependent prior density, it achieves exact probability matching. That is, $P_{\theta_{0}}\{\theta_{0}\leq\Pi^{-1}(\alpha\mid Y)\mid u\}=\alpha$ for all possible values of $\theta$ and $\alpha$. Then, by comparison with the constant prior, it follows that the latter produces posterior inferences which are biased towards lower values of $\theta$ under repeated sampling.
If the selection function is increasing, naive selective inference based on the unconditional model $Y\sim N(\theta,\sigma^{2})$ would lead to overestimation of $\theta$. Proposition 3 shows that correcting for selection via conditioning on the selection event has the opposite effect in a Bayesian setting: it leads to systematic underestimation of $\theta$. Note that for a realized $Y=y$, $C(y)=(-\infty,\Pi^{-1}(\alpha\mid y)]$ is a $\alpha 100\%$ credible interval for $\theta_{0}$, so the proposition states that all intervals of this form undercover the parameter. In García Rasines and Young (2022) it is shown that certain default priors, which depend on the sample size and assign low prior probability to parameters with low selection probability, can be used to correct this effect.
In view of Corollary 1, a similar behavior ought to arise in more general parametric settings provided the induced selection function on the normalized score, $p^{*}_{n}(z;\theta)$, is increasing. Figure 5 illustrates this phenomenon in the exponential model considered above, where all the induced selection functions are increasing. Qualitatively, we observe that when selection is deterministic there is a large deviation between the distribution of $\Pi(\theta_{0}\mid Y)$ and the uniform distribution, which would correspond with a perfectly calibrated posterior. This behavior persists from the Gaussian model to the exponential model. Randomizing selection reduces this discrepancy, but the deviation remains significant.
Figure 5: Cumulative distribution functions of $\Pi(\theta_{0}\mid Y^{n})$ in the exponential model (blue) with the same settings as before, and of the corresponding Gaussian posterior distributions (black). Left, $n=50$; right, $n=100$.

### v3

@@ 3.5 Inference on winners
where $Z_{n}^{1}=\sqrt{n}(\bar{Y}_{n}-\mu)$ and $Z_{n}^{2}=\sqrt{n}(V_{n}-\sigma^{2}+(\bar{Y}_{n}-\mu)^{2})$. The selection function is $p_{n}(y^{n})=\mathbf{1}(s>t)$, where $s$ is the $t$-value as defined above (using only the first $n_{1}$ observations).
If $n_{2}=0$, so that no further data is collected following selection, the asymptotic model for inference admits a tractable analysis. Upon rewriting the MLEs as a function of $(Z_{n}^{1},Z_{n}^{2})$, we obtain the asymptotic selection function
$p_{n}^{*}(z;\theta)=\mathbf{1}\left\{\sqrt{n}\mu+z_{1}>t\sqrt{\frac{z_{2}}{\sqrt{n}}+\sigma^{2}-\frac{z_{1}^{2}}{n}}\right\},$
which determines a selection event in the $(z_{1},z_{2})$-plane. If only $\mu$ is of interest, we can obtain the asymptotic marginal distribution of $Z_{n}^{1}$ by integrating out the variance component of the score:
$f(z_{1};\theta)\propto\phi\left(\frac{z_{1}}{\sigma}\right)\left[\Phi\left\{U(z_{1};\mu,\sigma^{2})\right\}-\Phi\left\{L(z_{1};\mu,\sigma^{2})\right\}\right],$
where
$L(z_{1};\mu,\sigma^{2})=\frac{z_{1}^{2}}{\sqrt{2}\sigma^{2}\sqrt{n}}-\frac{\sqrt{n}}{2},\quad U(z_{1};\mu,\sigma^{2})=\frac{\sqrt{n}}{\sqrt{2}\sigma^{2}}\left[\left(\frac{\sqrt{n}\mu+z_{1}}{t}\right)^{2}-\sigma^{2}+\frac{z_{1}^{2}}{n}\right].$
If $n_{2}>0$, the asymptotic selection function, $p_{n}^{*}(z;\theta)=E_{\theta}[p_{n}(Y^{n})\mid Z_{n}=z]$, does not admit, to the best of our knowledge, a simple analytical expression. It is the conditional acceptance probability of a subsample $t$-test given the full-sample sufficient statistics $(\bar{Y}_{n},V_{n})$. However, one may compute the asymptotic marginal density numerically:
$f(z_{1};\theta)\propto\phi\left(\frac{z_{1}}{\sigma}\right)p_{n}^{*}(z_{1};\theta),\quad p_{n}^{*}(z_{1};\theta)=\int_{-\infty}^{\infty}\phi\left(\frac{z_{2}}{2\sigma^{2}}\right)p^{*}_{n}(z_{1},z_{2};\theta)\mathrm{d}z_{2}.$
Figure 3 shows one-dimensional cuts of $p^{*}_{n}(z;\theta)$, $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r_{n}^{*}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ for $n_{1}=20$, $n_{2}=20$, $\theta=(0,1)$ and $t=1$. Under the true parameter, the selection probability is $\varphi_{n}(0,1)\approx 0.16$. The first two functions were computed via numerical integration, and $r_{n}^{*}(h;\theta)$ via Monte Carlo integration of $p^{*}_{n}(z;\theta)$. As expected, the approximation of the true selection probability by a Gaussian probability is much more accurate in the direction of the mean ($h_{1}$) than in the direction of the variance ($h_{2}$): in this problem $Z_{n}^{1}$ is exactly normally distributed. Additionally, since the selection condition is employed to determine large values of $\mu$, the selection function fluctuates much more in the first coordinate, corresponding to $\mu$ (top left plot), than in the second one (top right).
@@
In both cases, approximate selective $p$-values for $H_{0}\colon\mu\leq\mu_{0}$ vs. $H_{1}\colon\mu>\mu_{0}$ can be constructed from the asymptotic marginal distribution of $Z_{n}^{1}$ at $\mu=\mu_{0}$, where $\sigma^{2}$ is substituted by a consistent estimator:
$p_{\text{val}}(y^{n})=\int_{z_{1}}^{\infty}f(z_{1};\mu_{0},\hat{\sigma}^{2})\mathrm{d}z_{1},\quad z_{1}=\sqrt{n}(\bar{y}_{n}-\mu_{0}).$ (15)
Figure 4 shows the distribution of (15), estimated over 1,000 replications, for the case $n_{1}=100$, $n_{2}=50$ and test $H_{0}\colon\mu\leq 0$ vs. $H_{1}\colon\mu>0$. The first plot corresponds to the true generating value $\mu=0$, and the second one to $\mu=0.15$. We fixed $\sigma^{2}=1$ and $t=1$ in both cases. To produce the numerical results we have used the plug-in estimator
$\hat{\sigma}^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2},$
which is consistent as long as the selection probability does not vanish.
Figure 4: Repeated sampling distribution of $p$-values for the inference on winners model, under the null hypothesis (left), and an alternative hypothesis (right); $n_{1}=100$, $n_{2}=50$.
This example is considered further in Section 4.2, where it is seen that repeated sampling properties of Bayesian inference under the specified selection model can be well approximated by those in the asymptotically equivalent selective Gaussian location model.
@@ 4 Asymptotic posterior distributions
We elaborate further on the distinction between both constructions below.
@@
The growing popularity of selection-adjusted posteriors underscores the importance of understanding their theoretical properties, particularly, as we shall see, because their behaviour deviates from that of standard Bayesian models. This section makes two primary contributions towards such understanding. We prove that the consequences of Theorem 1 extend to the Bayesian framework, leading to a result related to the Bernstein von-Mises Theorem for the context of selection. Specifically, we show that the asymptotic shape of the posterior distribution matches that of the corresponding Gaussian selection model under a uniform prior. These latter models provide a more intuitive framework that can be leveraged to better understand the former. Furthermore, as a consequence of the previous result, we indicate a key practical limitation of Bayesian selective inference: probabilistic claims about the parameter (e.g. that it lies within a certain credible interval with probability 90%), cannot be given a frequentist interpretation with high accuracy. In order to achieve frequentist-matching properties, more complex prior distributions that depend on the sample size are required.
@@ 4.1 Asymptotic behavior of selective posteriors
$E_{\theta_{0}}(T_{n})\to 0\quad\text{and}\quad\sup_{\|\theta-\theta_{0}\|\geq\varepsilon}E_{\theta}(1-T_{n})\to 0.$ (20)
@@
where $Z_{n}=n^{-1/2}I_{\theta_{0}}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})\rvert_{\theta_{0}}$. That is, the Total Variation distance between the true and asymptotic posteriors vanishes asymptotically in probability conditionally on selection.
@@
4.2 Exponential model (ctd.)
Figure 5 compares the posterior densities for the local parameter $h=\sqrt{n}(\theta-\theta_{0})$, $\pi_{h,n}(h\mid y^{n})$, for two exponential models with their theoretical Gaussian limits. We examine two of the scenarios described at the end of Section 3: deterministic selection and randomization with Gaussian noise. The model is $Y_{1},\ldots,Y_{50}\sim\text{Exp}(\theta)$, $\theta\sim\text{Gamma}(1,0.1)$, and the true parameter is $\theta_{0}=2$. All posterior densities correspond to an observed sample average of $\bar{y}_{n}=0.45$. The true posterior densities based on the exponential likelihood, $\pi_{h,n}(h\mid y^{n})$, are shown in blue, while the Gaussian approximations $\pi^{*}_{h,n}(h\mid z_{n})$, as predicted by the theorem, are shown in black. For comparison, the posterior density in the absence of selection, which is approximately a rescaled $N(1/\bar{y}_{n},\theta_{0}^{2}/n)$ under standard theory, is shown in red. Two selection mechanisms are considered: a deterministic one (left), with selection condition $\bar{Y}_{n}<0.5$, and a randomized one (right), with selection condition $\bar{Y}_{n}+0.2\times N(0,1/n)<0.5$. In both cases, the selection probabilities $\varphi_{n}(\theta)$ and $\varphi^{*}_{n}(h)$ were obtained via numerical integration, as in the numerical investigation presented in Figure 2. We can readily observe that both posterior densities are non-Gaussian, exhibit higher uncertainty than the non-selective ones, have slowly decaying tails in the direction of the parameter space where $\varphi^{*}_{n}(h)$ vanishes, and demonstrate marked skewness, particularly in the case involving deterministic selection.
@@
4.3 Inference on winners (ctd.)
For further numerical illustration, we revisit the example discussed in Section 3.5, inference on a normal mean, under a selection condition defined by a $t$- test on a subsample. We assume true parameter value $\theta_{0}=(0,1)$ and consider sample sizes $n_{1}=100,n_{2}=50$, with threshold $t_{n}=1$, which corresponds to a selection probability of about 0.15. Figure 6 illustrates, for a particular data sample, the marginal posteriors for $h_{1}$ and $h_{2}$, derived from the exact posterior $\pi_{h,n}(h\mid y^{n})$, constructed assuming the prior $\pi(\mu,\sigma^{2})\propto 1/\sigma$, and the Gaussian approximation $\pi^{*}_{h,n}(h\mid z_{n})$. In each case, the full posterior for the local parameter $h=(h_{1},h_{2})$ is constructed by MCMC sampling. The vertical lines in the figure indicate the limits of 90% credible intervals for $h_{1},h_{2}$ constructed from the exact posterior $\pi_{h,n}(h\mid y^{n})$. In the case of $h_{1}$, for instance, this interval is (0.31,3.81)(-0.31,3.81) and has probability content under the marginal posterior corresponding to the Gaussian approximation of 0.902. The corresponding figure in the case of $h_{2}$ is 0.916. The marginal posteriors in the true selection model and the Gaussian approximation model match closely, but not exactly.
@@
$(20,20)$ 0 0.889 (0.058) 0.843 (0.076)
1 0.895 (0.059) 0.866 (0.066)
2 0.907 (0.055) 0.900 (0.051)
3 0.918 (0.064) 0.926 (0.069)
$(50,25)$ 0 0.897 (0.045) 0.865 (0.055)
1 0.910 (0.045) 0.886 (0.049)
2 0.918 (0.055) 0.909 (0.044)
3 0.920 (0.086) 0.919 (0.054)
$(100,50)$ 0 0.904 (0.028) 0.885 (0.040)
1 0.911 (0.028) 0.890 (0.038)
2 0.917 (0.030) 0.900 (0.037)
3 0.910 (0.076) 0.909 (0.039)
4.4 Frequentist mis-calibration of selective Bayesian inference
@@
In the classical IID framework, the Bernstein-von Mises Theorem establishes asymptotic equivalence between the Bayesian and frequentist approaches to inference on $\theta$. In practical terms, this implies that Bayesian posterior probabilities can be interpreted as frequentist probabilities under the assumption that there is a fixed, true data-generating parameter $\theta_{0}$. For instance, $1-\alpha$ Bayesian credible sets are also asymptotic $1-\alpha$ frequentist confidence sets. This equivalence stems from the fact that Gaussian location models with a uniform prior on the location satisfy this probability-matching condition exactly.
@@
In the presence of selection, this duality breaks: selection-adjusted posterior distributions can remain systematically mis-calibrated as the sample size increases. This follows by combining Proposition 1 with the observation that the uniform prior on the location parameter lacks a Bayesian-frequentist equivalence under selection, except in the trivial scenarios where the selective distribution is itself Gaussian (Proposition 2). Achieving approximate frequentist calibration in this setting typically requires the use of more carefully tuned priors that account for the selection mechanism and vary with the sample size: see García Rasines and Young (2022). This is not surprising, since, typically in selective inference, the dependence of the selection function on $n$ does not vanish asymptotically. This consideration of mis-calibration is relevant both for Bayesians and for frequentist practitioners who use Bayesian methods for their computational or interpretive convenience.
@@
Formally, a prior density for $\theta$ is (exactly) probability-matching if $\Pi(\theta_{0}\mid Y)\sim U(0,1)$ under repeated sampling of $Y$ when $\theta_{0}$ is the true parameter, where $\Pi(\cdot\mid Y)$ is the marginal posterior distribution function of $\theta$ induced by the said prior. This ensures that probabilistic claims about the parameter based on the posterior distribution admit a valid frequentist interpretation.
@@
Conveniently, uniform priors on location parameters are probability-matching. Thus, by the Bernstein-von Mises Theorem, any fixed prior density which is strictly positive at $\theta_{0}$ is first-order probability-matching, meaning that $\Pi(\theta_{0}\mid Y)\rightsquigarrow U(0,1)$ under $\theta_{0}$, and hence credible sets are asymptotically valid confidence sets.
By contrast, the following proposition shows that, except in the trivial cases, there is no fixed prior which is probability matching. This analysis encompasses valid improper priors, such as $\pi(\theta)\propto 1$, which arises as the prior density of the asymptotic posterior in Proposition 1. Consequently, it is often the case that the limiting Gaussian posteriors indicated in Proposition 1 are not calibrated in a frequentist sense. This proposition relates to a well-known one-dimensional result due to Lindley (1958), which states that the only exact probability-matching priors in one-dimensional models are uniform priors on location parameters. Throughout, for a symmetric matrix $B\in\mathbb{R}^{p\times p}$, we write $B\succ 0$ to mean that $B$ is positive definite, and $B\preceq 0$ to mean that $B$ is negative semidefinite.
@@
Let $Y$ follow a selective Gaussian distribution with base model $\{N(\theta,\Sigma)\colon\theta\in\mathbb{R}^{p}\}$ and selection function $p(y)$, where $\Sigma\succ 0$ is known. A probability-matching prior for $\theta$ exists if and only if the selection function can be written, almost everywhere, as
$\log p(y)=y^{\top}Ay+b^{\top}y+c\quad\text{for some }A\in\mathbb{R}^{p\times p},b\in\mathbb{R}^{p},c\in\mathbb{R},$
where $A=A^{\top}$ and $\Sigma^{-1}-2A\succ 0$. In this case, the selective distribution is
$Y\sim N\left((\Sigma^{-1}-2A)^{-1}(\Sigma^{-1}\theta+b),(\Sigma^{-1}-2A)^{-1}\right)$
and the probability matching prior is the improper uniform prior on $\theta$, $\pi(\theta)\propto 1$.
If $p$ is additionally required to be bounded, as is the case for the $p^{*}$ function in Theorem 1, then the exponential-quadratic representation must satisfy $A\preceq 0$ and $b\perp\ker(A)$, where $\ker(A)=\{v\in\mathbb{R}^{p}:Av=0\}$. In this case, the selection mechanism can be realized via subspace Gaussian randomization: there exist an integer $r\leq p$, a full row-rank matrix $M\in\mathbb{R}^{r\times p}$, a positive definite matrix $\Gamma\in\mathbb{R}^{r\times r}$, and a realization $u\in\mathbb{R}^{r}$ such that, with
$U=MY+W,\qquad W\sim N(0,\Gamma)\text{ independent of }Y,$
conditioning on $U=u$ induces the same exponential–quadratic selection function. In this case,
$\log p(y)=-\frac{1}{2}y^{\top}M^{\top}\Gamma^{-1}My+(M^{\top}\Gamma^{-1}u)^{\top}y+c\equiv y^{\top}Ay+b^{\top}y+c,$
where $A=-(1/2)M^{\top}\Gamma^{-1}M$ and $b=M^{\top}\Gamma^{-1}u$ (with $c$ constant given $u$). Additive Gaussian randomization is standard in selective inference; see García Rasines and Young (2023). Cases with $r<p$ correspond to randomization only in certain directions. For example, if $M=(1,0,\ldots,0)$, only the first coordinate of $Y$ is randomized.
The degree of departure from uniformity of the marginal posterior density depends on the strength of selection bias, that is, on how much the selection mechanism perturbs the distribution of the data under the true parameter. Miscalibration is typically more pronounced under deterministic selection, and less severe for smooth selection functions, such as those arising from carving and randomized selection mechanisms.
This lack of calibration can be made more precise in the one-dimensional setting when the selection function of the Gaussian model is monotonic. This scenario is of major practical interest as it encompasses selection mechanisms that favor large estimated effects. In such cases, the posterior distribution tends to overstate, under repeated sampling, regions of the parameter space with low selection probability, i.e. it “overcompensates” selection bias, as it leads to underestimation, as opposed to overestimation, of the true effect. This result extends Proposition 1 in García Rasines and Young (2022).
@@
$P_{\theta_{0}}\{\Pi(\theta_{0}\mid Y)\leq\alpha\mid u\}<\alpha\quad\forall(\alpha,\theta_{0})\in(0,1)\times\mathbb{R}.$ (22)
Uniform miscalibration across the parameter space underscores the inadequacy of the uniform prior for this type of selection. The proof of this result hinges on the existence of a function $\pi(\theta;y)$ which is strictly increasing in $\theta$ for every $y$ such that, if employed as a data-dependent prior density, it achieves exact probability matching. That is, $P_{\theta_{0}}\{\Pi(\theta_{0}\mid Y)\leq\alpha\mid u\}=\alpha$ for all possible values of $\theta$ and $\alpha$. Then, by comparison with the constant prior, it follows that the latter produces posterior inferences which are biased towards lower values of $\theta$ under repeated sampling.
If the selection function is increasing, the posterior distribution tends to place too much probability mass on small values of the parameter relative to what a well-calibrated posterior would. That is, while uncorrected selective inference based on the unconditional model $Y\sim N(\theta,\sigma^{2})$ would lead to overestimation of $\theta$, conditioning on the selection event has the opposite effect: it leads to systematic underestimation of $\theta$.
Note that, for a realized $Y=y$, $C(y)=(-\infty,\Pi^{-1}(\alpha\mid y)]$ is a $\alpha 100\%$ credible interval for $\theta_{0}$. The proposition therefore implies that credible intervals of this form undercover in the frequentist sense, i.e. have coverage strictly smaller than $\alpha$. In García Rasines and Young (2022) it is shown that certain default priors, which depend on the sample size and assign low prior probability to parameters with low selection probability, can be used to mitigate this effect. Conversely, the opposite effect occurs in the other direction: the credible intervals $C(y)=[\Pi^{-1}(\alpha\mid y),\infty)$ cover the parameter with probability larger than the nominal level of $1-\alpha$. In particular, it follows that $p_{\text{val}}(y)=\Pi(\theta_{0}\mid y)$ constitutes a valid, though conservative, one-sided $p$-value function for the test $H_{0}\colon\theta\leq\theta_{0}$, as $P_{\theta_{0}}\{p_{\text{val}}(Y)<\alpha\}<\alpha$.
In view of Proposition 1, a similar behavior ought to arise in more general parametric settings provided the induced selection function on the normalized score, $p^{*}_{n}(z;\theta)$, is increasing. Figure 7 illustrates this phenomenon in the exponential model considered above, where all the induced selection functions are increasing. Qualitatively, we observe that when selection is deterministic there is a large deviation between the distribution of $\Pi(\theta_{0}\mid Y)$ and the uniform distribution, which would correspond with a perfectly calibrated posterior. This behavior persists from the Gaussian model to the exponential model. Randomizing selection reduces this discrepancy, but the deviation remains significant.
Figure 7: Cumulative distribution functions of $\Pi(\theta_{0}\mid Y^{n})$ in the exponential model (blue) with the same settings as before, and of the corresponding Gaussian posterior distributions (black). Left, $n=50$; right, $n=100$. Dashed lines, deterministic selection; dotted lines, randomized selection.